OpenStudy (anonymous):

compared to the graph of y=x^2 where has the graph of y=x^2-3 moved?

OpenStudy (anonymous):

will give hand jube for answer

OpenStudy (anonymous):

@bubblegumpop<3

OpenStudy (anonymous):

@im.celibate

OpenStudy (anonymous):

Hello, my name is Celia, what do you need help with?

OpenStudy (anonymous):

uhhh the question i juzt posted....

OpenStudy (anonymous):

please and thank you hugs?

OpenStudy (anonymous):

Okay, so is this algebra, linear equations...?

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

OKay.. So retype the question, please.

OpenStudy (anonymous):

why its already written

OpenStudy (anonymous):

compared to the graph of y=x^2, where has the graph of y=x^2-3 moved?

OpenStudy (anonymous):

Where is the graph? I kind-of need that. And the first equation is y=x^2, and the second is y=x^2-3, do you see a difference?

OpenStudy (anonymous):

honestly idk im retarded idk how to graph it

OpenStudy (anonymous):

i have autism and downsyndrome

OpenStudy (anonymous):

SO this is the graph for equation #2

OpenStudy (anonymous):

Okay, @Twiztidsplifff , let's focus on the question.

OpenStudy (anonymous):

oki ill try so whats the answer? ill give u a hug and a taco.

OpenStudy (anonymous):

do you like transformers?

OpenStudy (anonymous):

Hmm. I have to work you through the question, not straight out give it to you. @TheSmartOne

OpenStudy (anonymous):

Do you think you could help him, I am not good at this xD

OpenStudy (anonymous):

cuz i like transformers :p

OpenStudy (anonymous):

smartone?

TheSmartOne (thesmartone):

You can plug them in to desmos.com/calulator and see the graphs...

TheSmartOne (thesmartone):

desmos.com/calculator

TheSmartOne (thesmartone):

I gtg I have a math test to take lol

OpenStudy (anonymous):

same bro this is my exam

OpenStudy (anonymous):

its -3 btw

OpenStudy (anonymous):

@TheSmartOne I need ur help at my question real fast...

OpenStudy (anonymous):

to thats addi bro wallahi

OpenStudy (anonymous):

What ?!? @Twiztidsplifff

OpenStudy (anonymous):

do worry about it

OpenStudy (anonymous):

The answer is -3.